OpenStudy (anonymous):

How do you solve this (m - 5)^3

OpenStudy (anonymous):

what doyou mean solve?

OpenStudy (anonymous):

that's an expression but you haven't said what you want to do with it

OpenStudy (anonymous):

How do you find the product

OpenStudy (anonymous):

Uh she means how do you find the product

OpenStudy (unklerhaukus):

\[(m - 5)^3=(m - 5)(m - 5)(m - 5)\]

OpenStudy (anonymous):

Yes but I don't get how you find the product

OpenStudy (unklerhaukus):

\[(m - 5)(m - 5)(m - 5)=m(m - 5)(m - 5)-5(m - 5)(m - 5)\]